OpenStudy (anonymous):

in sailing a fathom is 6 feet deep. How many yards are in 58 fathoms? (show work please)

OpenStudy (hunus):

How many feet are in a yard?

OpenStudy (anonymous):

then multiply that number by 58... since you got 58 fathoms

OpenStudy (dan815):

58 * 6 = 348 feet 348 feet = X yards

OpenStudy (dan815):

3 feet = 1 yard

OpenStudy (dan815):

348 / 3 = 116 yards

OpenStudy (dan815):

or simply from the start, A fathom = 6 feet or 2 Yards 58 fathoms is 2*58 = 116 yards

OpenStudy (anonymous):

@dan815 @iafnook @Hunus thank you! (: